40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 26 17:24:24 2017, Steve, Update, General, power outage powerOutage.png
    Reply  Thu Dec 28 17:22:02 2017, gautam, Update, General, power outage - CDS recovery 28.png
       Reply  Thu Jan 4 14:39:50 2018, gautam, Update, General, power outage - timing error 
          Reply  Fri Jan 5 21:54:28 2018, rana, Update, General, power outage - timing error huh.png
             Reply  Fri Jan 5 22:19:53 2018, gautam, Update, General, power outage - timing error 
                Reply  Sat Jan 6 18:27:37 2018, gautam, Update, General, power outage - IFO recovery EX_QPD_Quad1_Faulty.pdfTrillium_faulty.pdf
                   Reply  Mon Apr 16 22:09:53 2018, Kevin, Update, General, power outage - BLRM recovery AA.jpgADC.jpg
                      Reply  Tue Apr 17 09:57:09 2018, Steve, Update, General, seismometer interfaces BS_Tril_Intrf-1X5.jpgGurs_Intf-1X1.jpg
                         Reply  Wed Apr 18 20:33:19 2018, Kevin, Update, General, seismometer interfaces front.jpgback.jpgconnectors.jpg
                            Reply  Fri Apr 20 23:36:28 2018, Kevin, Update, PEM, Seismometer BLRMs front.jpgback.jpgdetail.jpg
                               Reply  Tue Apr 24 21:19:08 2018, Kevin, Update, PEM, Seismometer BLRMs modified_schematic.pdftop.jpgbottom.jpgBS_Seis_PSD.pdf
                                  Reply  Wed Apr 25 17:44:39 2018, Arnold, Update, PEM, PEM Anti-Alias wiring 
                                     Reply  Thu Apr 26 09:35:49 2018, Kevin, Update, PEM, PEM Anti-Alias wiring board.jpgSeismometerPSD.pdf
                                        Reply  Thu Apr 26 20:22:21 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections setup.pdfseismometerASD.pdf
                                           Reply  Wed May 16 21:02:22 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections ADC_noise.pdf
                                        Reply  Thu Jun 14 15:24:32 2018, Steve, Update, PEM, ADC DAC In Line Test Boards are in ADC_DAC_in_(1).JPG
    Reply  Tue Jan 2 15:43:35 2018, Steve, Update, VAC, pumpdown after power outage pumpdown_from_44_mTorr.png
       Reply  Wed Feb 7 15:50:42 2018, Steve, Update, VAC, IFO pressure monitoring CC_analog.png
Message ID: 13787     Entry time: Tue Apr 24 21:19:08 2018     In reply to: 13777     Reply to this: 13788
Author: Kevin 
Type: Update 
Category: PEM 
Subject: Seismometer BLRMs 

In the ongoing attempt to recover the seismometer BLRMS, I removed the AA board from the rack and modified the BS seismometer Z channel. The BS_Z BLRMs seem to be recovered after this modification.

I removed the three resistors from the output of the circuit and wired the input and from the seismometer directly to the input to the ADC. The modified schematic is shown in attachment 1. Attachments 2 and 3 show the top and bottom of the modified board. The board is doing nothing now other than serving as a connector for this channel.

I put the board back in the rack and injected a 2 Vpp signal into the BS_Z channel and saw +/- 1600 cts in C1PEM-SEIS_BS_Z. I then plugged the seismometer back into the board and took the spectrum shown in attachment 4. This shows the working Z channel giving a reasonable seismic spectrum. Note that X and Y are not modified yet.

If there are no objections, I will modify all the other channels on the board in the same way tomorrow.

Attachment 1: modified_schematic.pdf  45 kB  Uploaded Tue Apr 24 22:35:47 2018  | Hide | Hide all | Show all
modified_schematic.pdf
Attachment 2: top.jpg  1.168 MB  Uploaded Tue Apr 24 22:36:00 2018  | Show | Hide all | Show all
Attachment 3: bottom.jpg  453 kB  Uploaded Tue Apr 24 22:36:11 2018  | Show | Hide all | Show all
Attachment 4: BS_Seis_PSD.pdf  38 kB  Uploaded Tue Apr 24 22:36:21 2018  | Hide | Hide all | Show all
BS_Seis_PSD.pdf
ELOG V3.1.3-